Part II: Show your work and explain your reasoning Chapter 8 F. Three gliders of masses 4.00 kg, 10.0kg, and 3.00 kg move on a frictionless horizontal track with x-velocities of 5.00 m/s, 3.00 m/s, and -4.00 m/s, as shown. Velcro couplers make the gliders stick together after colliding. 5.00 m/s 3.00 m/s -4.00 m/s (a) Find the final velocity of the train of three gliders. [Answer: 2.241 m/s] (b) Does your answer require that all the gliders collide and stick together at the same time? What if they collide in a different order? Explain your answers.
Answer #1

(a) Momentum conservation

Pi = Pf

(m1u1 + m2u2 + m3u3) = (m1 + m2 +m3) v

((4 x 5) + (10 x 3) - (3 x 4)) = (4 + 10 + 3) v

38 = 17 v

v = 2.235 i m/s (Answer)

(b) Yes according to the part (a) they both will collide and stick at same time thats why the initial momentum of all three is separate and final momentum is with all three as one body.

If they collide in a different time momentum conservation will be applied for 2 times
1st for the gliders who collide initially and then they stick together and assuming as one body
2nd for the 3rd glider and these 2 block body.